# A Scalable Quantum Non-local Neural Network for Image Classification

### v1

@@ Introduction
Figure 3: The pipeline for the Hybrid Classical-QNL-Net Models used in this work: (a) CNN-QNL-Net: Two convolutional layers with ReLU activation and max pooling extract features from input images, followed by a dropout layer and flattening of the features, and then two fully connected (FC) layers to prepare the data for the QNL-Net module. (b) PCA-QNL-Net: PCA reduces the input data dimensionality to 4 components, which are then processed by a fully connected (FC) layer before being fed into the QNL-Net module. (c) QNL-Net: The quantum part of the pipeline which performs computation on the four features obtained from the classical models according to the mechanism and outputs a single measurement obtained from qubit 0. (d) Post-QNL-Net Classical Computation: A fully connected layer is used to fine-tune the quantum output, and then this output is transformed and concatenated as desired for binary classification.
It has been shown in recent years that the integration of classical and quantum mechanisms for machine learning is very promising as these hybrid models leverage the strengths of both paradigms. Classical models perform best with feature extraction and dimensionality reduction problems, while quantum models offer superior capabilities by using quantum mechanical phenomena such as entanglement and superposition, enabling the extraction of exponentially more information out of each chunk of data (Biamonte et al. 2017; Huang et al. 2022). Therefore, this hybrid approach can lead to more effective machine learning systems, particularly for tasks involving pattern recognition in large and complex datasets.
@@
In this work, we employ two hybrid classical-quantum non-local neural network (QNL-Net) models: CNN-QNL-Net and PCA-QNL-Net, as seen in Fig. 3. The primary purpose for using these classical models is for dimensionality and feature reduction, transforming input data into a feature vector of size $4\times 1$, such that each value of the vector can be encoded into one qubit of the QNL-Net layer to capture and analyze long-range dependencies and intricate patterns among these features.
After measuring the output from the QNL-Net, we perform further classical computation by adding a fully connected layer consisting of a single learnable parameter, which helps fine-tune and optimize the quantum output and improve the model’s performance. The linear transformation performed by this layer is as follows.
@@
where $Q\in\mathbb{R}^{1}$ is the single QNL-Net output, $W_{4}\in\mathbb{R}^{1\times 1}$ is the weight matrix, and $b_{4}\in\mathbb{R}^{1}$ is the bias vector. Finally, for performing binary classification, we compute the complement of the probabilities and concatenate them together to obtain a single vector compatible for loss computation and backpropagation,
$\hat{y}=cat(x,1-x),$ (16)
where $\hat{y}\in\mathbb{R}^{2}$ is the resultant vector containing the predicted probabilities for both classes.
Quantum Simulation
@@
In our experiments, we utilize two widely used image processing datasets for image classification: MNIST (Deng 2012) and CIFAR-10 (Krizhevsky, Hinton et al. 2009) (sample images are shown in the Appendix in Fig. 5). For MNIST, we focus on binary classification using the digits 0 and 1, comprising 12,665 training samples and 2,115 testing samples. For CIFAR-10, we perform binary classification using classes 2 (birds) and 8 (ships). Before feeding the images into the models, we normalize them using the global mean and standard deviation of each dataset, scaling pixel values from [0, 255] to [0, 1].
@@
The experiments are conducted on a MacBook Pro with an M2 Max chip and 64GB RAM. The QNL-Net network is implemented by utilizing the EstimatorQNN module of Qiskit Machine Learning 0.7.2 and Qiskit 1.1.0, which facilitates the encoding of classical data into quantum data and enables the training of the ansatz. The classical node is constructed, and gradient optimization is performed using PyTorch 2.3.0, which connects seamlessly with the EstimatorQNN module. The models are trained for 100 epochs with a batch size of 1 using the negative log-likelihood (NLL) loss function for convergence. The Adam optimizer (Kingma and Ba 2014) is configured with different learning rates, which vary between 0.0001 to 0.0004 depending upon the model and the ansatz used as listed in Appendix Table 1. The ExponentialLR scheduler (Li and Arora 2019) uses a $\gamma$ decay rate of 0.9. The rest of the parameters are set to default.
@@
The experiments were conducted using different combinations of feature map repetitions $r$ = 1, 2, or 3 and the number of ansatz repetitions $D$ = 1, 2 or 3. The accuracies reported are averaged for all runs for each specific ansatz and model configuration, as shown in Appendix Table 1.
The results on the MNIST dataset for classes 0 and 1 indicate that the CNN-QNL-Net model performs slightly better than the PCA-QNL-Net model, achieving a near-perfect average classification test accuracy of 99.96% whereas the PCA-QNL-Net achieved a test accuracy of 99.59%. Ansatz-0 and Ansatz-2 generally yield better results compared to Ansatz-1 for this dataset, as evident in Appendix Table 1. For the CIFAR-10 dataset, the hybrid-QNL-Net models perform comparatively worse than the MNIST due to the introduction of three color channels (i.e., RGB) compared to MNIST’s grayscale images. However, the CNN-QNL-Net was still able to obtain an average test accuracy of 93.98%. Ansatz-1 performs better on the testing dataset for CIFAR-10 compared to the other ansatzes. On both datasets, the CNN-QNL-Net significantly outperforms the PCA-QNL-Net due to its ability to efficiently extract features from the dataset before feeding them to the QNL-Net module.
Increasing the depth of both the feature map and the ansatzes (i.e., $r$ and $D$, respectively) generally improved classification accuracies due to the increased expressiveness of the circuit. However, it also resulted in longer training times compared to using fewer repetitions, which still obtained reasonably good results within the bounds reported in Appendix Table 1. Furthermore, PCA-QNL-Net models required higher learning rates to obtain convergence than the CNN-QNL-Net models. The PCA-QNL-Net demonstrated faster training compared to the CNN-QNL-Net, as CNNs add an overhead for training parameters, utilizing a total of 34,282 classical parameters on MNIST and 41,314 classical parameters on CIFAR-10. The PCA-QNL-Net, however, optimizes only 22 classical parameters from the linear layers, offering an advantage in terms of classical training efficiency.
@@
Table 1 presents the complete results for all different ansatz and models run on the datasets. Further, Table 2 compares the results from our models to the benchmark quantum binary classifiers discussed in the Related Work. Our hybrid classical-quantum models, particularly the CNN-QNL-Net, outperform the benchmark models across both datasets. This superior performance underscores the effectiveness of combining classical convolutional networks with quantum neural network layers, enabling more robust and accurate classification.
Table 1: This table summarizes the performance metrics of various configurations of the QNL-Net model on binary classification tasks across datasets: MNIST digits 0 and 1 and CIFAR-10 classes 2 (bird) and 8 (ship). The configurations vary by the ansatz type, the model used (CNN-QNL-Net or PCA-QNL-Net), and the learning rate. The reported metrics include the average training accuracy and test accuracy across different variations of the number of repetitions of the feature map ($r$) and the number of repetitions of the ansatz ($D$). Numbers following $\pm$ represent the standard deviation across configurations.
@@
Table 2: This table compares the performance of our QNL-Net model with other benchmark quantum classifiers on binary classification tasks using the MNIST dataset. The CNN-QNL-Net model demonstrates better results using significantly fewer qubits.
@@
QTN-VQC (Qi, Yang, and Chen 2023) 0, 1 12 98.6
Hybrid TTN-MERA (Grant et al. 2018) 0, 1 8 $99.87\pm 0.02$
@@
SQNN (Wu, Tao, and Li 2022) 3, 6 64 97.47
QF-hNet-BN (Jiang, Xiong, and Shi 2021) 3, 6 12 98.27
@@
Figure 4: Training loss convergence and accuracy plots for CNN-QNL-Net and PCA-QNL-Net models using three ansatzes with one feature map repetition ($r=1$) and one ansatz repetition ($D=1$). (a) and (b) show loss convergence during training for the MNIST dataset (classes 0 and 1) and the CIFAR-10 dataset (classes 2 and 8), respectively. (c) and (d) display training accuracy on the respective datasets.
For the MNIST dataset, Figure 4a shows that the PCA-QNL-Net model achieves a lower final loss compared to the CNN-QNL-Net model across all ansatz configurations, suggesting that the PCA-QNL-Net model fits the training data more closely by the end of the training period. Despite this, the CNN-QNL-Net model reaches higher training accuracy more quickly than the PCA-QNL-Net model, as shown in Figure 4c, indicating better generalization. In contrast, on the CIFAR-10 dataset, the CNN-QNL-Net model demonstrates a more rapid decrease in training loss and achieves a lower final loss value compared to the PCA-QNL-Net model (Fig. 4b). This suggests faster convergence and a better fit to the CIFAR-10 training data. The CNN-QNL-Net model also consistently achieves higher training accuracy faster than the PCA-QNL-Net model in Fig. 4d, reinforcing its effectiveness in learning from more complex data. Overall, our experimental results confirm that integrating quantum circuits with classical neural network architectures significantly enhances model performance, setting a new benchmark in quantum machine learning.
@@ Discussion
The novelty of the QNL-Net architecture lies in its efficient utilization of fewer qubits, a critical consideration in the NISQ era. However, its significance extends beyond that to take into account the fundamental principles of quantum mechanics, particularly in its treatment of rotations around axes and entanglement. TThe choice of rotation gates in the QNL-Net ansatzes is tied to the fundamental idea behind non-local neural networks, which aim to capture intricate spatial dependencies within the data. In the quantum paradigm, rotation gates achieve this by translating quantum states around different axes, thereby implementing spatial transformations.
Now, for simplification, consider the raw data $\mathrm{x}$ and the three embeddings ($\theta,\phi,g$) in the non-local neural net architecture as akin to the four features passed on to the QNL-Net circuit. In non-local neural nets, raw data is not embedded, but in this work, we utilize rotation around the $z$-axis primarily because it allows us to maintain a trainable parameter on qubit 0 after the feature map encoding which ensures the consistency and completeness of quantum theory while not affecting the measurement initially. The $R_{y}$ gates around qubits 1 and 2 for embeddings $\theta$ and $\phi$, respectively, is crucial. These gates induce phase shifts in the states of the corresponding qubits, playing a vital role in generating interference effects and complex quantum states. Additionally, they facilitate capturing relative phase information, which is pivotal in discerning subtle patterns within the data. The linear embedding $g$ translates to qubit 3, where an $R_{x}$ gate modifies the probability amplitudes of states |$\ket{0}$ and |$\ket{1}$, fine-tuning the data representation within the quantum circuit.
@@
A key advantage of QNL-Net is leveraging quantum entanglement. While classical non-local blocks typically involve matrix multiplications within the embeddings and element-wise summation with the raw data, the QNL-Net utilizes the intrinsic significance of CNOT entanglements to replicate variable dependencies present in classical non-local mechanisms. This creates a highly interconnected system between all qubits, effectively extracting intricate probabilities and facilitating the exploration of complex data structures in a quantum framework. As we can infer from the results, it does not matter significantly what the specific pattern for entanglement is in different ansatz, rather the idea that all qubits encompassing different rotations are entangled together. Then, further applying another $R_{z}$ gate on qubit 0, which is already in a highly entangled state, introduces additional phase modulation within the entangled system. This modulation is crucial for fine-tuning the quantum state before measurement, thereby influencing the outcome probabilities.
@@
Despite the promising results obtained from these models, several limitations were identified. The current implementation is restricted by the reliance on classical computing methods that might be computationally inefficient as the datasets become much more extensive and the models are more complex. Still, it is a trade-off we must consider as we utilize lesser quantum resources. Multi-class classification posed a particular challenge as it performed poorly, likely due to the small circuit size with Pauli-$Z$ measurement at only one qubit, which limits the model’s capability in distinguishing between multiple classes. Additionally, the model’s performance could be further validated by testing on more diverse and larger datasets and exploring the effects of different quantum encodings and variational circuit designs.
@@
Nevertheless, the implications of this research extend to various practical applications. In fields like image classification, medical imaging, and real-time video analysis, efficiently capturing long-range dependencies using quantum-enhanced models can lead to significant advancements in accuracy and performance. Furthermore, the scalable nature of QNL-Net suggests that as quantum hardware evolves, these models could be deployed in real-world scenarios, providing a competitive edge over traditional classical approaches.
@@ Conclusion
This paper introduced the Quantum Non-local Neural Networks (QNL-Net) mechanism as a novel hybrid classical-quantum approach for image classification. Through experiments on MNIST and CIFAR-10 datasets, QNL-Net models demonstrated competitive performance in binary classification tasks, using fewer qubits compared to traditional quantum classifiers. The use of fundamental quantum techniques like entanglement and rotation gates proved effective in capturing intricate spatial dependencies critical for image analysis.
However, QNL-Net exhibits limitations in multi-class classification and efficiency with larger, complex datasets due to reliance on classical preprocessing methods. These challenges underscore the need to explore future work on innovative QNL-Net variants and optimization techniques. Additionally, exploring the integration of more efficient quantum encoding strategies might also enhance performance.
In conclusion, QNL-Net promises advancements in accuracy and efficiency for image classification tasks, with potential transformative impacts in fields requiring robust pattern recognition, such as medical imaging and real-time video analysis. Its ability to operate with reduced computational resources compared to classical and existing quantum methods positions QNL-Net as a scalable solution for quantum-enhanced machine learning applications, laying the groundwork for broader use in practical applications.
Supplementary Material: Technical Appendix
Background on Quantum Computing
Quantum computing leverages the fundamentals of quantum mechanics, such as superposition and entanglement, to introduce new properties to computing. It uses the concept of qubits that have the computational basis states |$\ket{0}$ and |$\ket{1}$, which can also be represented as
@@
A qubit is a linear combination of these basis states, which is the general principle of superposition, and can be represented as a vector in a two-dimensional complex Hilbert space, such that,
@@
The first step to any quantum computation is encoding classical data into quantum states. To achieve this, several fundamental encoding techniques are used: (i) Basis encoding maps classical bits 0 and 1 to states |$\ket{0}$ and |$\ket{1}$ directly, and therefore each classical bit string is encoded as the corresponding quantum state; (ii) Amplitude encoding uses the amplitudes of a quantum state to represent classical data such that the sum of the squared amplitudes of the quantum state is normalized to 1 for the classical data; (iii) Phase encoding maps classical information to the phases of a quantum state and is used in the proposed QNL-Net mechanism.
Quantum computations are performed primarily by manipulating quantum states through unitary transformations, achieved using quantum gates. Hadamard (H) gate is used to attain an equal superposition of the two basis states. The H gate maps the basis state |$\ket{0}$ to |$\frac{\ket{0}+\ket{1}}{\sqrt{2}}$ and the basis state |$\ket{1}$ to |$\frac{\ket{0}-\ket{1}}{\sqrt{2}}$. Rotation gates ($R_{x},R_{y},R_{z}$) rotate the state of a qubit around a specified axis on the Bloch sphere. A Phase ($P$) gate shifts the phase of a qubit by a specified angle $\phi$ such that applying $P(\phi)$ to |$\ket{\psi}$ in eq.(18) results in $P(\phi)\ket{\psi}=\alpha\ket{0}+\beta e^{i\phi}\ket{1}$. A CNOT (CX gate) is a two-qubit gate that flips the state of the second qubit (target) only if the first qubit (control) is |$\ket{1}$. The following are the matrix representations of the relevant gates utilized in this work:
$H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\hskip 5.69054ptP(\phi)=\begin{bmatrix}1&0\\ 0&e^{i\phi}\end{bmatrix},$ (19)
[…]
$R_{y}(\theta)=\begin{bmatrix}cos(\frac{\theta}{2})&-sin(\frac{\theta}{2})\\ sin(\frac{\theta}{2})&cos(\frac{\theta}{2})\end{bmatrix},$ (21)
$R_{z}(\theta)=\begin{bmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{bmatrix},\hskip 5.69054ptCX=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{bmatrix}.$ (22)
CNN-QNL-Net
@@
In the proposed CNN-QNL-Net architecture, we start with two convolutional layers, each with an activation function and max pooling, for an input image tensor $X\in\mathbb{R}^{W\times H\times C}$, where $W$ is the width, $H$ is the height, and $C$ is the number of channels (i.e. 1 for grayscale images and 3 for RGB images) of the input image. In general, mathematically, a convolution operation ‘*’ for an input image $I$ and a filter $K$ to output a feature map $F$ looks like,
@@
where $i,j$ are positions in the output feature map $F$, $m,n$ are positions in the filter $K$ for channel $c$, and $M$, $N$, and $C$ are the width, height, and number of channels of the filter respectively. The first convolutional layer applies a 2D convolution operation with $K_{1}$ filters (or kernels) of size $5\times 5$ resulting in $K_{1}$ output channels, and is defined as,
@@
Then, we flatten the output $D$ to obtain a 1-dimensional vector $F$ of size $(W_{4}\cdot H_{4}\cdot K_{2})\times 1$. We apply a fully-connected (FC) layer to this flattened vector $F$ with ReLU activation, which results in,
@@
where $\mu$ is the calculated mean vector of the data, $1_{N}$ is an N-dimensional vector of ones, and $\bar{X}$ is the centered input data matrix. We perform SVD on this centered matrix to decompose it into several component matrices with various interesting properties (Brunton and Kutz 2022),
@@
where $\mu_{z}$ and $\sigma_{z}$ are the mean and standard deviation of $Z$ respectively. $\bar{Z}$ is a reduced vector of size $L\times 1$ (i.e. $4\times 1$) and is then passed to a fully connected layer, such that,

### v2

@@ Introduction
Figure 3: The proposed hybrid Classical-quantum QNL-Net frameworks comprises the following components: (a) CNN-QNL-Net: Two convolutional layers with ReLU activation and max-grouping are used to extract features from input images. Subsequently, a dropout layer is applied, followed by flattening of the features. Two fully connected (FC) layers are utilized to preprocess the data for integration with the QNL-Net. (b) PCA-QNL-Net: Principal Component Analysis (PCA) is utilized to reduce the dimensionality of the input data to 4 components. These components are processed by a fully connected (FC) layer before being input into the proposed QNL-Net. (c) QNL-Net: This component of the pipeline conducts computations on the four features obtained from the classical models within the mechanism and generates a single measurement from qubit 0. (d) Post-QNL-Net Classical Computation: A fully connected layer is employed to fine-tune the quantum output. Subsequently, this output is transformed and concatenated as required for binary classification.
@@
In this study, we employ two hybrid classical-quantum QNL-Net models: CNN-QNL-Net and PCA-QNL-Net. Classic models are used for dimensionality and feature reduction, converting input data into a feature vector of size $4\times 1$. Each value in this vector corresponds to one qubit in our QNL-Net layer, facilitating the capture and analysis of long-range dependencies and intricate patterns among these features. The incorporation of CNN architecture into our QNL-Net Module is driven by CNN’s ability to capture spatial dependencies and identify local patterns within complex image data through convolutional and pooling layers, effectively reducing input dimensionality while retaining essential features.
In the proposed CNN-QNL-Net architecture, the process begins with two convolutional layers, each integrating an activation function and max pooling, to process an input image tensor $X\in\mathbb{R}^{W\times H\times C}$. Here, $W$ represents the width, $H$ denotes the height, and $C$ signifies the number of channels in the input image (e.g., 1 for grayscale images and 3 for RGB images). The input data matrix is $X\in\mathbb{R}^{N\times P}$, where $N$ is the number of samples in the dataset and $P$ is the total number of pixels per image. The linear transformation performed by this layer is as follows.
@@
where $Q\in\mathbb{R}^{1}$ is the single QNL-Net output, $W_{4}\in\mathbb{R}^{1\times 1}$ is the weight matrix, and $b_{4}\in\mathbb{R}^{1}$ is the bias vector. Further details on the CNN and PCA architectures are given in Appendix. Our model uses the negative log-likelihood (NLL) loss function for the binary classification problem. The NLL loss measures the variation between the true labels $y$ and the classical predicted probabilities $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]$ obtained from the measurement of the hybrid quantum-classical QNL-Net model, and is defined for binary classification as:
$\mathcal{L}(\phi)=-\sum^{N}_{i=1}\left(y_{i}\log\hat{y}_{i1}+(1-y_{i})\log\hat{y}_{i0}\right),$ (18)
where $N$ is the number of samples in the dataset, $\hat{y}_{0}$ is the predicted probability for class 0, $\hat{y}_{1}$ is the predicted probability for class 1, and $\phi$ are the classical parameters. The gradient calculations are described in the Appendix.
Evaluation Results
@@
In our Qiskit simulations, we used two widely used image processing datasets for image classification: MNIST (Lecun et al. 1998) and CIFAR-10 (Krizhevsky, Sutskever, and Hinton 2012) (sample images are shown in Appendix in Fig. 5). For MNIST, we focus on binary classification using the digits 00 and $1$, comprising $12,665$ training samples and $2,115$ testing samples. For CIFAR-10, we perform a binary classification using classes $2$ (birds) and $8$ (ships). Before feeding the images into the models, we normalize them using the global mean and standard deviation of each data set, scaling the pixel values from [$0,255$] to [$0,1$].
@@
In our study, we extensively evaluated the performance of the proposed QNL-Net in benchmark datasets, MNIST and CIFAR-10, demonstrating its resistance to noise and potential advantages over traditional quantum binary classification models. The Qiskit simulations are performed on a MacBook Pro computer equipped with an M2 Max chip and $64$GB RAM. Our QNL-Net has been implemented using the EstimatorQNN module of Qiskit Machine Learning $0.7.2$ and Qiskit $1.1.0$ (Javadi-Abhari et al. 2024), allowing the encoding of classical data into quantum data and facilitating the training of the ansatz. The Qiskit ZFeatureMap is utilized to leverage the quantum-enhanced feature space, providing a quantum advantage in classification tasks (Havlíček et al. 2019). Following the measurement of our QNL-Net output, further classical computation is performed by adding a fully connected layer with a single learnable parameter to fine-tune and optimize the quantum output, thereby enhancing the model’s performance. The classical node is constructed and gradient optimization has been performed using PyTorch $2.3.0$ (Paszke et al. 2017), seamlessly integrated with the EstimatorQNN module. The models are trained for $100$ epochs with a batch size of $1$, using the negative log-likelihood (NLL) loss function for convergence. The Adam optimizer (Kingma and Ba 2014) is configured with varying learning rates between $0.0001$ and $0.0004$, depending on the model and ansatz used, as detailed in Appendix Table 1. The ExponentialLR scheduler (Li and Arora 2019) with a decay rate of $0.9$ is used, while the remaining parameters were set to default.
@@
Simulations are performed using all combinations of the number of repetitions for the encoder ($r$ = $1,2$ or $3$) and the number of repetitions for the VQC ($D$ = $1,2$ or $3$), with results averaged across the nine possible configurations. The averaged accuracies across all runs for each specific ansatz and model configuration are given in Table 1. The learning rates are found by grid search, and given in Table 1 for these settings. Fig. 4 illustrate the models’ performance and convergence behavior during training for $D=r=1$. The simulation results obtained from the MNIST dataset for classes 00 and $1$ reveal that our CNN-QNL-Net model marginally outperforms our PCA-QNL-Net model, achieving an average classification test accuracy of $99.96\%$. In contrast, PCA-QNL-Net achieved a test precision of $99.59\%$. Ansatz-0 and Ansatz-2 generally exhibit superior performance compared to Ansatz-1 for this dataset, as indicated in Table 1. However, for the CIFAR-10 dataset, hybrid QNL-Net models demonstrate relatively lower performance compared to MNIST due to the inclusion of three color channels (RGB) as opposed to the grayscale images in MNIST. Nonetheless, our CNN-QNL-Net model still achieves an average test accuracy of $93.98\%$. Ansatz-1 performs better on the testing dataset for CIFAR-10 compared to the other ansatzes. Across both datasets, our CNN-QNL-Net model significantly outperforms our PCA-QNL-Net model due to its efficient feature extraction capabilities before inputting the data into the QNL-Net architecture.
Figure 4: Training loss convergence and test accuracy plots for the CNN-QNL-Net and PCA-QNL-Net models for the three ansatzes with one feature map repetition ($r=1$) and one ansatz repetition ($D=1$). (a) and (b) display the training loss convergence on MNIST and CIFAR-10, respectively. (c) and (d) show the corresponding test accuracy on these datasets.
We note that our proposed PCA-QNL-Net models required higher learning rates compared to our CNN-QNL-Net models. The proposed PCA-QNL-Net exhibited faster training in contrast to our CNN-QNL-Net, as CNNs entail an overhead for training parameters, utilizing a total of $34,282$ classical parameters for MNIST and $41,314$ classical parameters for CIFAR-10. Further, there are $5D+4r$ classical parameters in the quantum gates and $22$ classical parameters for the linear layers. Details of the gradient optimization calculations are provided in the Appendix.
@@
Table 2 compares the outcomes from our models to the benchmark quantum binary classifiers: QTN-VQC (Qi, Yang, and Chen 2023), Hybrid TTN-MERA (Grant et al. 2018), Tensor Ring VQC (Peddireddy et al. 2022), SQNN (Wu, Tao, and Li 2022), and QF-hNet-BN (Jiang, Xiong, and Shi 2021), where QTN-VQC, Hybrid TTN-MERA, and Tensor Ring VQC have results in their papers on classes 0 and 1 in MNIST; and SQNN and QF-hNet-BN have results for classes 3 and 6 in their papers. Our hybrid classical-quantum models, particularly the proposed CNN-QNL-Net, outperforms the benchmark models across both datasets. This superior performance underscores the effectiveness of integrating classical convolutional networks with quantum neural network layers, enabling more robust and accurate classification.
Table 1: Performance of the proposed QNL-Net model on binary classification tasks across datasets: MNIST digits 00 and $1$ and CIFAR-10 classes $2$ (bird) and $8$ (ship).
@@
Table 2: Performance of the QNL-Net model compared with QTN-VQC (Qi, Yang, and Chen 2023), Hybrid TTN-MERA (Grant et al. 2018), Tensor Ring VQC (Peddireddy et al. 2022), SQNN (Wu, Tao, and Li 2022), and QF-hNet-BN (Jiang, Xiong, and Shi 2021) on binary classification tasks using the MNIST dataset. The CNN-QNL-Net model demonstrates better results using significantly fewer qubits.
@@
QTN-VQC $0,1$ 12 $98.60$
Hybrid TTN-MERA $0,1$ 8 $99.87\pm 0.02$
Tensor Ring VQC $0,1$ 4 $99.30$
@@
SQNN $3,6$ $64$ $97.47$
QF-hNet-BN $3,6$ $12$ $98.27$
@@
For the MNIST dataset, the results depicted in Figure 4 indicate that the proposed PCA-QNL-Net model achieves a lower final loss compared to our proposed CNN-QNL-Net model across all ansatz configurations. This suggests that the PCA-QNL-Net model closely fits the training data by the end of the training period. However, the CNN-QNL-Net model attains a higher test accuracy more rapidly than the PCA-QNL-Net model, as illustrated in Figure 4, indicating a superior generalization. In contrast, with respect to the CIFAR-10 data set, the CNN-QNL-Net model exhibits a faster reduction in training loss and achieves a lower final loss value compared to the PCA-QNL-Net model (Fig. 4b). This implies faster convergence and better fit to the CIFAR-10 training data. Furthermore, the CNN-QNL-Net model consistently achieves higher test accuracy with more speed than our PCA-QNL-Net model in Fig. 4, reinforcing its efficacy in learning from complex data.
Overall, our experimental findings validate that the integration of quantum circuits with classical neural network architectures significantly enhances the performance of the QNL-Net model, establishing a new benchmark in quantum machine learning.
@@ Discussion
The proposed QNL-Net leverages quantum entanglement as a key advantage. Although classical non-local blocks typically involve matrix multiplications within the embeddings and element-wise summation with the raw data, the QNL-Net utilizes $C_{X}$ entanglements to replicate variable dependencies present in classical non-local mechanisms. This fosters a highly interconnected system among all qubits, effectively extracting intricate probabilities and exploring complex data structures within a quantum framework. The specific pattern for entanglement in different ansatzes is not as crucial as the entanglement of all qubits with different rotations, creating a highly entangled state. Furthermore, applying another $R_{z}$ gate to the qubit, 00, introduces phase modulation within the entangled system, crucial for fine-tuning the quantum state prior to measurement, thus influencing the probability of an outcome.
@@
Despite the promising results obtained from these models, several limitations were identified. The current implementation is restricted by the reliance on classical computing methods that might be computationally inefficient as the datasets become much more extensive and the models are more complex. Still, it is a trade-off we must consider as we utilize lesser quantum resources. Multi-class classification posed a particular challenge as it performed poorly, likely due to the small circuit size with Pauli-$Z$ measurement at only one qubit, which limits the model’s capability in distinguishing between multiple classes. Furthermore, the performance of the model could be further validated by testing on larger and more diverse datasets and exploring the effects of different quantum encodings and variational circuit designs.
@@
Nevertheless, the implications of this research extend to various practical applications. In fields like image classification, medical imaging, and real-time video analysis, efficiently capturing long-range dependencies using quantum-enhanced models can lead to significant advancements in accuracy and performance. Furthermore, the scalable nature of our QNL-Net suggests that, as quantum hardware evolves, these models could be deployed in real-world scenarios, providing a competitive edge over traditional classical approaches.
@@ Conclusion
This work introduces a Quantum Non-local Neural Networks (QNL-Net) framework as a novel hybrid classical-quantum approach for image classification. Through experiments on MNIST and CIFAR-10 datasets, our proposed QNL-Net models demonstrate competitive performance in binary classification tasks, utilizing fewer qubits compared to traditional quantum classifiers. The incorporation of fundamental quantum entanglement and rotation gates proves to be effective in capturing intricate spatial dependencies crucial for image analysis.
Appendix
Introduction to Quantum Computing
Quantum computing utilizes the principles of quantum mechanics, such as superposition and entanglement, to enhance traditional computing capabilities. At the core of quantum computing are qubits, the quantum counterparts to classical bits. These qubits can simultaneously exist in states of |$\ket{0}$ and |$\ket{1}$ due to superposition, providing a significant increase in computational power compared to classical bits (Preskill 2018), can be represented as
@@
A qubit embodies a linear combination of the computational basis states, illustrating the fundamental concept of superposition. This state can be succinctly depicted as a vector within a two-dimensional complex Hilbert space, characterized by the following mathematical expression:
@@
Quantum computations are performed primarily by manipulating quantum states through unitary transformations, achieved using quantum gates. Hadamard (H) gate is used to attain an equal superposition of the two basis states. The H gate maps the basis state |$\ket{0}$ to |$\frac{\ket{0}+\ket{1}}{\sqrt{2}}$ and the basis state |$\ket{1}$ to |$\frac{\ket{0}-\ket{1}}{\sqrt{2}}$. Rotation gates ($R_{x},R_{y},R_{z}$) rotate the state of a qubit around a specified axis on the Bloch sphere. A Phase ($P$) gate shifts the phase of a qubit by a specified angle $\lambda$, such that, applying $P(\lambda)$ to |$\ket{\psi}$ in eq.(20) results in
$P(\lambda)\ket{\psi}=\alpha\ket{0}+\beta e^{i\lambda}\ket{1}.$ (21)
A CNOT ($C_{X}$ gate) is a two-qubit gate that flips the state of the second qubit (target) only if the first qubit (control) is |$\ket{1}$. The following are the matrix representations of the relevant gates utilized in this work:
$H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\hskip 5.69054ptP(\lambda)=\begin{bmatrix}1&0\\ 0&e^{i\lambda}\end{bmatrix},$ (22)
$R_{x}(\lambda)=\begin{bmatrix}cos(\frac{\lambda}{2})&-i\hskip 2.84526ptsin(\frac{\lambda}{2})\\ -i\hskip 2.84526ptsin(\frac{\lambda}{2})&cos(\frac{\lambda}{2})\end{bmatrix},$ (23)
$R_{y}(\lambda)=\begin{bmatrix}cos(\frac{\lambda}{2})&-sin(\frac{\lambda}{2})\\ sin(\frac{\lambda}{2})&cos(\frac{\lambda}{2})\end{bmatrix},$ (24)
$R_{z}(\lambda)=\begin{bmatrix}e^{-i\frac{\lambda}{2}}&0\\ 0&e^{i\frac{\lambda}{2}}\end{bmatrix},\hskip 5.69054ptC_{X}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{bmatrix}.$ (25)
CNN Architecture
@@
In the proposed CNN-QNL-Net architecture, we start with two convolutional layers, each with an activation function and max pooling, for an input image tensor $X\in\mathbb{R}^{W\times H\times C}$, where $W$ is the width, $H$ is the height, and $C$ is the number of channels (i.e., 1 for grayscale images and 3 for RGB images) of the input image. In general, mathematically, a convolution operation ‘*’ for an input image $I$ and a filter $K$ to output a feature map $F$ looks like,
@@
where $i,j$ are positions in the output feature map $F$, and $m,n$ are positions in the filter $K$ for channel $c$. $M$, $N$, and $C$ are the width, height, and number of channels of the filter, respectively. The first convolutional layer applies a 2D convolution operation with $K_{1}$ filters (or kernels) of size $5\times 5$ resulting in $K_{1}$ output channels, and is defined as,
@@
Further, we flatten the output $D$ to obtain a 1-dimensional vector $F$ of size $(W_{4}\cdot H_{4}\cdot K_{2})\times 1$. We apply a fully-connected (FC) layer to this flattened vector $F$ with ReLU activation, which results in,
@@
where $\mu$ is the calculated mean vector of the data, $1_{N}$ is an N-dimensional vector of ones, and $\bar{X}$ is the centered input data matrix. We perform SVD on this centered matrix to decompose it into several component matrices,
@@
where $\mu_{z}$ and $\sigma_{z}$ are the mean and standard deviation of $Z$ respectively. $\bar{Z}$ is a reduced vector of size $L\times 1$ (i.e., $4\times 1$) and is then passed to a fully connected layer, such that,